ANALYSIS OF MICROWAVE SPECTRUM, INTERNAL ROTATION AND C–H⋯F INTERACTIONS OF THE CHF_{3} ⋯C_{2}H_{3}F WEAKLY BOUND COMPLEX

\underline{LENA F. ELMUTI}, DANIEL A. OBENCHAIN, REBECCA A. PEEBLES, SEAN A. PEEBLES, Department of Chemistry, Eastern Illinois University, 600 Lincoln Ave., Charleston, IL 61920.

C–H⋯X hydrogen bonds with systematic variation of halogens (X = F, Cl, Br) have been examined using Fourier-Transform Microwave (FTMW) spectroscopy. Rotational constants for trifluoromethane–vinyl fluoride (TFM⋯VF) were consistent with a Cs symmetry structure that exhibited both bifurcated and single C–H⋯F interactions between the TFM and VF. This near prolate asymmetric top exhibited three-fold internal rotation of the CF_{3} group causing characteristic doubling in its spectra. Initial assignments were completed using chirped-pulse FTMW spectroscopy with additional measurements made using a resonant-cavity FTMW spectrometer. Rotational constants from ab initio calculations at the MP2/6-311++G(2d,2p) level were in agreement with preliminary experimental values (A = 4828 MHz, B = 1049 MHz, C = 1018 MHz). XIAM^{a} was used to provide the barrier to internal rotation (25(5) cm^{−1}) and other spectral information. Spectroscopic parameters for the normal isotopic species as well as preliminary structural results on the C–H⋯F interactions in this complex will be presented.

aH. Hartwig and H. Dreizler, Z. Naturforsch, \underline{51a}, (1996), 923-932.